OpenStudy (anonymous):

the residents of a city voted on whether to raise property taxes,the ratio of yes votes to no votes was 8 to 5.if there were 4175 no votes,what was the total number of votes?

OpenStudy (anonymous):

e n e

OpenStudy (anonymous):

Number of yes votes/Number of NO votes =8/5 Number of yes votes/4175=8/5

OpenStudy (anonymous):

Now, can u find number of yes votes?

OpenStudy (anonymous):

And FIND TOTAL VOTES = NUMBER of yes votes + NUMBER of NO votes